The product of two consecutive integers is 18 more than the square of the smaller number. Find the integers.... I can't figure out how to use algebra in order to solve this. Thank you for your help in advance!
Found 2 solutions by richard1234, mananth:
Answer by richard1234(7193)

Let x, x+1 be the integers. Then , so the integers are 18 and 19.

Answer by mananth(16949)

let the integers be x and (x+1)
x(x+1) = x^2+18
x^2+x=x^2+18
x=18
so the integers are 18,19